### v3

# Moments and the Basic Equilibrium Equations

Abstract.
We develop the classical theory of moments of configurations of weighted points with a focus on systems with an identically vanishing first moment. The latter condition produces equations for equilibrium configurations of systems of interacting particles under the sole condition that interactions are pairwise and along the line determined by each pair of particles. Complying external forces are admitted, so the description of some dynamical equilibrium configurations, such as relative equilibria in Celestial Mechanics, is included in our approach. Moments provide a unified framework for equilibrium problems in arbitrary dimensions. The equilibrium equations are homogeneous, of relatively low degree, invariant by oriented isometries (for interactions depending only on mutual distances), and are obtained through simple algebraic procedures requiring neither reduction by isometries nor a variational principle for their determination. Our equations include the renowned sets of $n$-body central configuration equations by O. Dziobek, and by A. Albouy and A. Chenciner. These equations are extended to a rather broad class of equilibrium problems, and several new sets of equilibrium equations are introduced. We also apply moments to establish a theory of constraints for mutual distances of configurations of fixed dimension and of co-spherical configurations, thus re-obtaining and adding to classical results by A. Cayley, among others.
Key words and phrases:
Affine Geometry, Barycentric Coordinates, Equilibria, $N$-Body Problem, Central Configurations.
2020 Mathematics Subject Classification:
70F10, 37N05, 70Fxx, 37Cxx.
1. Introduction
Equilibrium problems have contributed to the development of Mathematics and Physics since ancient Greece. The work of Archimedes on the principle of the lever is considered the beginning of the mathematical treatment of Statics, but even before him, there had appeared books attributed to Aristotle and to Euclid on the dynamics and balance of the lever, respectively [15]. The concept of barycenter, or center of gravity, of a system of point masses, and its application to studying equilibrium configurations go back to Archimedes. As we shall review in section 2, the barycenter is defined for sets of weighted points whose total weight, $\mu_{0}$, is nonzero as the unique point where the first moment, $\mu_{1}$, vanishes. In the 17th century, C. Huygens used $\mu_{1}$ to introduce the notion of expected value in probability, and realized the importance of the second moment, $\mu_{2}$, in the form of the moment of inertia (Euler’s terminology), for the analysis of pendulum motion. Later on G. W. Leibniz and J. S. König applied $\mu_{2}$ in Geometry and Dynamics [19]. A consequential event was the invention by A. F. Möbius of homogeneous, barycentric coordinates in the early 19th century [13]. In contemporary Mathematical Statistics and Analysis, among other areas, one finds the classical (Stieltjes, Hausdorff, Hamburger) moment problem: given a sequence of moments of a real random variable, determine the (unique) probability distribution (measure) that produced the sequence [27, 30].
In the physical sciences, it has been a long tradition to represent configurations of point masses, electric charges, point vortices, etc, by a finite set of points $X$ in the Euclidean affine space $\mathbb{A}=\mathbb{R}^{N}$ over the field of scalars $K=\mathbb{R}$, and a function $w$ from $X$ to $K$ denoting the mass, charge, vorticity, etc, of each point of $X$. In an abstract sense, we refer to $X$ and $w$ as configuration and weight function, respectively, and call the pair $(X,w)$ a weighted system. Three elementary functions are associated with every weighted system: the total weight $\mu_{0}$ (“zeroth moment”), the first moment $\mu_{1}$ and the second moment $\mu_{2}$. In the case of a configuration of point masses, $\mu_{0}$ is the total mass and, besides the previously mentioned direct relation between the barycenter and $\mu_{1}$, and between the moment of inertia and $\mu_{2}$, if one allows the configuration to vary differentiably with time, then $\mu_{1}$ can be used to define the total linear momentum, while $\mu_{2}$ becomes (twice) the kinetic energy. All these concepts naturally extend to discrete and continuous settings.
Drawing apart from tradition, in this paper we view the interactions of particles as defining the weight functions themselves. More precisely, we introduce a family of weighted systems $(X,w_{x})$, $x\in X$, associated with a configuration $X$ of interacting particles, so that $w_{x}(y)$ is the force (this familiar term will be often used in place of interaction) exerted by the particle $y\in X$ on the particle $x$, while $w_{x}(x)$ is the negative of the total force acting on $x$. This basic remark, and the key realization that the equilibrium configurations correspond to fulfilling the condition $\mu_{1}=\overrightarrow{O}$ (this notation means that the first moment is the null function) for each system $(X,w_{x})$, motivates a deeper investigation of the theory of moments of weighted systems in Euclidean affine spaces. At the same time, we have a bridge to applications of moments to problems of equilibrium configurations of interacting particles possibly subject to certain types of external forces. To be more explicit, all our equilibrium equations are valid for systems of interacting weighted points satisfying the single fundamental hypothesis:
(FH): The interaction between the particle at $x$ and the particle at $y$, with $x,y\in X$, is directed along the line determined by $x$ and $y$.
As a matter of fact, our formalism holds for systems satisfying the weaker hypothesis:
(WFH): The total interaction (e.g., internal plus external forces) at each point $x$ of the configuration $X$ is a linear combination of the relative position vectors from $x$ to the remaining points of $X$.
Hypothesis (WFH) extends the scope of applications of this work to systems affected by some inertial forces, for instance, centrifugal forces which give rise to relative equilibrium configurations.
We shall see that the theory of moments of weighted systems provides both a unified framework for the construction of families of equilibrium equations with distinctive properties, as well as a tool to study geometric aspects of the configurations themselves. Two important points of departure for our work were the classical identities deduced by the pioneers of moment theory, and the investigations on central configurations of the $n$-body problem.
Central configurations is a research topic with a long history and recognized relevance to Celestial Mechanics [4, 3, 28, 31, 34]. One of the reasons is that the only known explicit solutions of the Newtonian $n$-body problem, namely the homographic motions, start at such configurations (cf. Wintner [34], §375). The possible homographic motions include the relative equilibria, which are planar periodic solutions, and, in recent times, have been shown to occur in any Euclidean space of even dimension [3, 28]. In the context of the $n$-body problem, it is common to use the terminologies planar (or two-dimensional) central configurations and relative equilibria interchangeably. Notwithstanding the many particular questions that have been answered, so far a classification of central configurations is known only for three arbitrary masses, thanks to Euler [17] and Lagrange [26], and for four equal masses, thanks to Albouy [2], who based his work on a previously obtained symmetry result by himself [1]. General, basic problems on central configurations remain unsolved. An outstanding example is the celebrated Smale’s sixth problem for the 21st century [32, 33], which asks whether the number of similarity classes of relative equilibria of $n$ positive point masses is finite for all $n$. In the broader sense of central configurations of arbitrary dimension, Smale’s problem is part of the Chazy-Wintner-Smale finiteness conjecture [4]. The finiteness problem has been fully solved only for $n=4$, by Hampton and Moeckel [22]. Considering just generic finiteness with respect to the masses, finiteness has been proved for $n=5$ in dimension two by Albouy and Kaloshin [5], and for the central configurations of $n$ bodies in dimension n2n-2, the so-called Dziobek configurations, by Moeckel [29] (see also Dias [14]). Below we consider the most important sets of central configuration equations used in the classification and finiteness problems, interspersed with summaries of our present contributions.
In the seminal paper [16], O. Dziobek used mutual distances as coordinates to describe central configurations as critical points of a modified potential function subject to constraints that guaranteed the configuration was planar. The constraints used by Dziobek were expressed in terms of the determinants introduced by A. Cayley [10], which later became known as Cayley-Menger determinants [8, 9]. For the planar four-body problem, Dziobek produced relatively simple algebraic equations expressing the reciprocal cubed mutual distances $r_{ij}^{-3}$ as an affine function of the product of suitably defined (volume) barycentric coordinates, $\Delta_{i}\Delta_{j}$, divided by the product of the masses, $m_{i}m_{j}$. His approach was adopted and developed by Albouy in the previously mentioned classification of four-equal-mass central configurations [1, 2]. Later on, Moeckel [29] adapted Dziobek’s method to dimension n2n-2, and, using barycentric coordinates, Albouy [4] provided a simple algebraic path to the Moeckel-Dziobek equations which did not require the (typically large) Cayley-Menger determinant.
Notably the whole Dziobek formalism is contained in the theory of weighted systems $(X,w)$ with $\mu_{1}=\overrightarrow{O}$. The weight functions of such systems (with fixed $X$) form a vector space, $\mathbb{W}_{0}(X)$, and we describe a method for constructing bases for $\mathbb{W}_{0}(X)$ using volume barycentric coordinates. The case $\dim\mathbb{W}_{0}(X)=1$ of our theory, and the fundamental hypothesis (FC) of page 1, are implicit in Albouy’s algebraic approach: in essence, it is the ideas in reference [4] that we recast and generalize to $\dim\mathbb{W}_{0}(X)\geq 1$. Besides the intensive use of barycentric coordinates, this generalization requires some additional concepts, in particular that of codimension of a configuration, which plays a chief role throughout the remainder of this paper. A significant number of equilibrium (including central configuration) equations is introduced through the “codimension-exterior power formula”
(1.1) $\bigwedge^{c+1}\mathbb{W}_{0}(X)=\{O\},$
where c=nd1c=n-d-1 is the codimension of $X$, $n$ is the number of points in $X$, and $d$ is dimension of the affine closure of $X$, usually referred to as the dimension of $X$. It would not be an overstatement to say that (1.1) synthesizes the basic theory of systems with $\mu_{1}=\overrightarrow{O}$. Formula (1.1) is used in section 5 to provide a broad generalization of some determinantal identities obtained by Dziobek [16, 21] for central configurations with codimension one. These identities yield tailored equilibrium equations that exploit properties such as symmetry and dimension, as illustrated in example 5.3. In view of its versatility, we deem equation (1.1) one of the main contributions of our work. We also remark that the Dziobek-type homogeneous, quadratic central configuration equations of Dias [14] are here extended to arbitrary weighted systems with $\mu_{1}=\overrightarrow{O}$, and we show that they actually provide a characterization of such systems having any specific dimension. Finally, a somewhat surprising connection with Algebraic Geometry arises as we interpret barycentric coordinates as Plücker coordinates: Plücker relations [24, 25] for Grassmannians of $d$-dimensional subspaces of $\mathbb{R}^{n-1}$ emerge directly from the generalization of Dias’ formulæ.
Regarding applications, two remarks seem in order. Firstly, Plücker relations, the linear relations for volume barycentric coordinates which form our basis vectors of $\mathbb{W}_{0}(X)$, and the generalized Dziobek equations of section 6 may be used to formulate equilibrium problems entirely in terms of barycentric coordinates, via corollary 2.6. And secondly, all the weight functions defined by the models of interacting particles having $X$ as an equilibrium configuration coinhabit the space $\mathbb{W}_{0}(X)$. Thus they are a fortiori interrelated, and their relationships may be expressed through equations such as those produced by formula (1.1). In this context, our bases for $\mathbb{W}_{0}(X)$ supply a purely geometric (i.e. non-physical) common framework for representing the weight functions of each model individually.
Going back to the topic of central configurations, it is worth noting that Dziobek equations also found an application in the pioneering resolution of the finiteness problem for gravitational relative equilibrium configurations [22], as well as for stationary configurations of vortices [23]. However, the main formulation of these finiteness problems employed a different set of central configuration equations, namely the one introduced by A. Albouy and A. Chenciner in the celebrated article [3]. Briefly, the finiteness proofs in [22, 23] applied ideas from Algebraic Geometry to exclude the possibility of a continuum of solutions, taking advantage of the fact that the so-called Albouy-Chenciner equations are algebraic. A remarkable feature of the Albouy-Chenciner equations is that they depend only on the mutual distances of the configuration which, besides eliminating the rotational and translational symmetries, makes them at once valid for configurations of all dimensions.
Unlike the known deductions of the Albouy-Chenciner equations [3, 22], which use reduction by isometries, early on in the paper we obtain these equations for every system with $\mu_{1}=\overrightarrow{O}$ as a nearly straightforward consequence of the basic interplay of the moments $\mu_{0}$ and $\mu_{1}$ with the second moment $\mu_{2}$. Later on, we provide another set of algebraic equations exhibiting the properties of the Albouy-Chenciner equations mentioned in the previous paragraph. We call these new equations the extended Leibniz identities, as they are closely related to a well-known identity involving moments and mutual distances attributed to Leibniz. We stress that, just as for the generalized Dziobek approach, the applications of the generalized Albouy-Chenciner and the extended Leibniz identities include all equilibrium problems satisfying the weaker fundamental hypothesis (WFC) stated on page 1. Moreover, concerning Geometry, the interplay of the first three moments allows us to establish a theory of constraints for distances from the points of a configuration to an arbitrary point in the ambient space, and for mutual distances between the points of a configuration of prescribed dimension or lying on a sphere. In particular, Cayley-Menger determinantal conditions for the dimension are derived, and the number of independent Cayley-Menger constraints needed to ensure the configuration has a given dimension is expressed in terms of the configuration’s codimension $c$ simply as $c+1\choose 2$}. Our formula for the number of constraints generalizes the ones for dimensions 1, 2 and 3 in Wintner [34], $\S 357$. It underscores once again the importance of the codimension in the study of configurations of points in Euclidean affine spaces.
Before moving on to a summary of specific contents, we make a brief note related to the moment problem mentioned in the first paragraph. We notice that the moment problem suggests an inverse of the equilibrium problem treated in the present work, viz., assuming the configuration $X$ is known, one would be asked to find the interactions $w_{x}$, $x\in X$, which make $X$ an equilibrium. For this problem, our equilibrium equations seem to offer relatively simple conditions to determine the $w_{x}$. In particular, the interactions can be determined using the linear systems provided by equations (2.2) or (4.3). If the interactions are symmetric, i.e., if Newton’s third law of motion (popularly “Action == (1)(-1) $\times$ Reaction”) holds, then, using the generalized Dziobek equations of section 6, the $w_{x}$ can be directly expressed in terms of barycentric coordinates (or mutual distances) and $c+1\choose 2$} parameters, where $c$ is the codimension of $X$.
Outline of Specific Contents
The fundamentals of the theory of moments of weighted systems are reviewed in section 2, culminating with the Huygens-Leibniz-König (HLK) theorem, which collects classical results relating $\mu_{0},\mu_{1}$ and $\mu_{2}$. We demonstrate three corollaries of the HLK theorem with conditions on the second moment $\mu_{2}$ of a weighted system with $\mu_{0}=0$ to have $\mu_{1}=\overrightarrow{O}$. Prototypes of the generalized Albouy-Chenciner equations and the extended Leibinz identities appear at this point. Together with barycentric coordinates, which are introduced in section 4, the HLK theorem and its corollaries are the main tools for our formulations of equilibrium problems.
Equilibrium equations for systems of interacting particles are firstly discussed in section 3. According to our fundamental hypothesis (FC), the total force on each particle $x\in X$ is a linear combinations with coefficients $\varphi_{x,y}$ of the relative positions from $x$ to the remaining particles $y\in X\setminus\{x\}$. We apply the results of section 2 to obtain equilibrium equations that are either linear or bilinear on the $\varphi_{x,y}$, and linear on the squared mutual distances between points of $X$. These equations contain, in particular, the generalization of the Albouy-Chenciner equations and the extended Leibniz identities.
Section 4 is dedicated to the set of all weighted systems with a fixed configuration $X$ and $\mu_{1}=\overrightarrow{O}$. The corresponding weight functions form a vector space $\mathbb{W}_{0}(X)$ whose dimension is the codimension of $X$. Equation (1.1) immediately follows. In section 4, we view Dziobek configurations as simplexes $S$ with an additional point in the affine closure of $S$. This geometric interpretation allow us to introduce devices called Dziobek trees for $X$ in order to construct explicit bases for $\mathbb{W}_{0}(X)$. Section 4 closes with a set of homogeneous second degree equations that characterize the weight functions in $\mathbb{W}_{0}(X)$. Remark 4.10 discusses applications to Grassmannians.
Section 5 concerns some illustrative applications of the previous sections to central configurations. After providing a formal definition of central configurations followed by a short review of basic results, some classical formulæ  are reobtained, and new central configuration equations are discussed. We consider the not so explored problem of central configurations with zero total mass, for which some geometric properties and algebraic relations are directly determined. In addition, we examine the planar five-body central configurations having three collinear bodies, assuming the total mass is nonzero. Our results impose some significant restrictions on the possible central configurations which are consistent with previous studies of particular cases by other authors.
Section 6 focuses on generalizing the set of equations found by Dziobek [16] to describe central configurations as constrained critical points. Dziobek’s equations assume their simplest form when $c=1$, and, in section 6, using matrices, we extend Albouy’s arguments [4] to any symmetric family of $n$ weight vectors in $\mathbb{W}_{0}(X)$ on a common configuration $X$ of arbitrary dimension. For systems of interacting particles, symmetry means the interactions obey Newton’s third law of motion, which, in the notation of section 3, translates simply as $\varphi_{x,y}=\varphi_{y,x}$, for all $x,y\in X$.
The article concludes with two appendices. In appendix A we show how the HLK theorem of section 2 can be used to find constraints for mutual distances. We apply corollary 2.6 to obtain constraints corresponding to either fixing the dimension of the configuration $X$, which take the form of vanishing Cayley-Menger determinants of Dziobek subconfigurations of $X$, or imposing co-sphericity, which are expressed as the vanishing of the determinant of the so-called relative configuration matrix [3, 28] of $X$. Two theorems are proved in appendix A. In theorem A.2.1 we clarify the relationship between the space of weight functions with $\mu_{1}=\overrightarrow{O}$, $\mathbb{W}_{0}(X)$, and the kernels of the relative configuration matrix and the Cayley-Menger matrix of $X$. The dimension of the configuration is thus related to vanishing Cayley-Menger determinants of subconfigurations. Then, in theorem A.2.2, we obtain the expression $c+1\choose 2$} for the number of independent constraints that ensures a configuration has a fixed dimension. Finally, in appendix B, we determine the kernel of the relative configuration matrix viewed, as in reference [3], as a bilinear form. The extended Leibniz identities follow therefrom, and a final set of central configuration formulæ  is provided.
2. Moments of a Weighted System
We begin with a brief overview of affine geometry; see [8] for a thorough exposition. Let $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$ be an Euclidean affine space, that is, a triple formed by a point set $\mathbb{A}$, a vector space $\overrightarrow{\mathbb{A}}$ over a field $K$ endowed with an inner product $\cdot$ , and a free action $\tau$ of the additive group of $\overrightarrow{\mathbb{A}}$ on $\mathbb{A}$. The vector $\overrightarrow{v}\in\overrightarrow{\mathbb{A}}$ acts on the point $x\in\mathbb{A}$ by translating it to a point $y\in\mathbb{A}$ and is unique in that sense, so we write
$\overrightarrow{xy}=\overrightarrow{v}.$
The zero vector, $\overrightarrow{O}$, corresponds to the trivial translation. For any pair of points $x,y\in\mathbb{A}$, it is assumed the existence of a vector $\overrightarrow{v}\in\overrightarrow{\mathbb{A}}$ satisfying the above equation; in other words, $\tau$ is transitive. It is common to write
$x+\overrightarrow{xy}=y\quad\text{or}\quad\overrightarrow{xy}=y-x.$
The pair $(\overrightarrow{\mathbb{A}},\tau)$ is the affine structure of the affine space $(\mathbb{A},\overrightarrow{\mathbb{A}},\tau)$, which is often denoted simply by $\mathbb{A}$. A basic and very useful result is the Chasles relation: $\overrightarrow{xy}+\overrightarrow{yz}=\overrightarrow{xz}$.
Let $X$ be a finite subset of an Euclidean affine space $\mathbb{A}$ and $w$ be a function from $X$ to $K$, the field of scalars of $\mathbb{A}$. We refer to $X$ as a configuration of points in $\mathbb{A}$, and to $w$ as a weight function on $X$. Three key maps can be assigned to the weighted system formed by $X$ and $w$, i.e., to the pair $(X,w)$.
Definition 2.1.
The total weight $\mu_{0}$, the first moment $\mu_{1}:\mathbb{A}\longrightarrow\overrightarrow{\mathbb{A}}$ and the second moment $\mu_{2}:\mathbb{A}\longrightarrow K$ of the weighted system $(X,w)$ are respectively given by
$\mu_{0}=\sum_{x\in X}w(x),\ \ \ \mu_{1}(p)=\sum_{x\in X}w(x)\overrightarrow{px},\ \ \text{and}\ \ \mu_{2}(p)=\sum_{x\in X}w(x)\overrightarrow{px}^{2},$
where the square of a vector stands for the inner product of the vector with itself.
The total weight can be thought of as the “zeroth moment”, hence its notation. It is sometimes convenient to see $\mu_{0},\mu_{1}$ and $\mu_{2}$ as (linear) maps on the weight function.
We state some basic results regarding the moments $\mu_{0}$, $\mu_{1}$ and $\mu_{2}$ of a fixed weighted system $(X,w)$ in $\mathbb{A}$.
Lemma 2.2.
Given $p,q\in\mathbb{A}$, we have that $\mu_{1}(p)-\mu_{1}(q)=\mu_{0}\overrightarrow{pq}.$
The proof is straightforward.
If $\mu_{0}\neq 0$, there is a unique point $G\in\mathbb{A}$ such that $\mu_{1}(G)=\overrightarrow{O}$, namely
$G=q+\frac{1}{\mu_{0}}\sum_{x\in X}w(x)\overrightarrow{qx},$
where $q$ is an arbitrary point of $\mathbb{A}$. We call $G$ the barycenter of $(X,w)$ and denote it by $\text{bar}(X,w)$. Using the barycenter makes it possible to simplify the expressions of the first and second moments, as shown by proposition 2.3 and the Huygens-Leibniz-König theorem below.
Proposition 2.3.
Let $p\in\mathbb{A},x\in X$.
  1. (1)
    If $\mu_{0}\neq 0$ and $G=\text{\emph{bar}}(X,w)$, then $\mu_{1}(p)=\mu_{0}\overrightarrow{pG}.$
  2. (2)
    If $\mu_{0}=0$ and $w(x)\neq 0$, then $\mu_{1}(p)=w(x)\overrightarrow{G_{x}x}$, where $G_{x}$ is the barycenter of the weighted system $\left(X\setminus\{x\},w\big|_{X\setminus\{x\}}\right)$.
Proof.
Part (1) follows directly from lemma 2.2 and the definition of $G$. For part (2), note that $\mu_{1}$ is constant according to lemma 2.2, so we have that $\mu_{1}(p)=\mu_{1}(x)$. But the total weight of the weighted system $\left(X\setminus\{x\},w\big|_{X\setminus\{x\}}\right)$ is $w(x)\neq 0$ and its first moment at the point $x$ is $\mu_{1}(x)$. So, by applying part (1) to the latter system with $p=x$, we obtain part (2). ∎
Part (2) of proposition 2.3 has a close relation to barycentric coordinates, see remark 4.4. We have the following immediate consequence of lemma 2.2 and proposition 2.3.
Corollary 2.4.
  • (a)
    $\mu_{0}\neq 0$ if and only if $\mu_{1}$ is a bijection.
  • (b)
    $\mu_{0}=0$ if and only if $\mu_{1}$ is constant.
In particular, $\mu_{1}=\overrightarrow{O}$ implies $\mu_{0}=0$.
Regarding the second moment and its relation to the total weight and first moment, we have the following key theorem to which, for briefness, we shall refer by the acronym HLK.
Theorem (Huygens-Leibniz-König).
For every $p,q\in\mathbb{A}$, we have that
(2.1) $\mu_{2}(p)-\mu_{2}(q)=\overrightarrow{pq}\cdot(\mu_{1}(p)+\mu_{1}(q)).$
Therefore
  1. (1)
    if $\mu_{0}\neq 0$ and $G=\text{\emph{bar}}(X,w)$, then $\mu_{2}(p)-\mu_{2}(G)=\mu_{0}\,\overrightarrow{pG}^{2}$, for all $p\in\mathbb{A}$,
  2. (2)
    if $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{v}$, then $\mu_{2}(p)-\mu_{2}(q)=2\overrightarrow{v}\cdot\overrightarrow{qp},$ for all $p,q\in\mathbb{A}$, and
  3. (3)
    for every $q\in\mathbb{A}$, the Leibniz’s identity
    $\mu_{0}\mu_{2}(q)-\mu_{1}(q)^{2}=\sum_{\{x,y\}\subset X}w(x)w(y)\overrightarrow{xy}^{2}$
    holds.
Proof.
In order to obtain equation (2.1), we just have to multiply both sides of the elementary identity
$\overrightarrow{px}^{2}=\overrightarrow{qx}^{2}+\overrightarrow{pq}^{2}+2\overrightarrow{qx}\cdot\overrightarrow{pq}$
by $w(x)$, add for all $x\in X$, and apply lemma 2.2. Picking $q=G$ gives (1), and taking $q\in\mathbb{A}$ arbitrary produces (2). To prove (3), set $p=y\in X$. From (2.1), we can write
$\sum_{y\in X}w(y)\mu_{2}(y)=\mu_{0}\mu_{2}(q)-\mu_{1}(q)^{2}+\sum_{y\in X}w(y)\,\overrightarrow{yq}\cdot\mu_{1}(y).$
By lemma 2.2, we can replace $\mu_{1}(y)$ by $\mu_{0}\overrightarrow{yq}+\mu_{1}(q)$ in the latter summation. Leibniz’s identity follows.
The HLK theorem has a broad range of applications. Below we state just two of its corollaries, the second of which is consequential to a significant part of this work, including the study of configurations described in terms of mutual distances, see appendix A.
Corollary 2.5.
We have that $\mu_{1}=\overrightarrow{O}$ if and only if $\mu_{0}=0$ and
$\sum_{\{x,y\}\subset X}w(x)w(y)\overrightarrow{xy}^{2}=0$
In appendix B we discuss an extension of corollary 2.5.
Corollary 2.6.
Suppose $\mu_{0}=0$. The following assertions are equivalent:
  • (i)
    $\mu_{1}=\overrightarrow{O}$;
  • (ii)
    $\mu_{2}$ is constant;
  • (iii)
    the restriction of $\mu_{2}$ to $X$ is constant.
Proof.
If $\mu_{1}=\overrightarrow{O}$, then $\mu_{2}$ is constant according to part (2) of the HLK theorem. Thus we have (i) $\Rightarrow$ (ii); (ii) $\Rightarrow$ (iii) is immediate. If the restriction of $\mu_{2}$ to $X$ is constant and $\mu_{1}=\overrightarrow{v}$ then, from the HLK theorem, part (2),
$0=\mu_{2}(y)-\mu_{2}(x)=2\overrightarrow{v}\cdot\overrightarrow{xy},\quad\text{for all}\ x,y\in X.$
But, for all $y\in X$,
$\overrightarrow{v}=\mu_{1}(y)=\sum_{x\in X}w(x)\overrightarrow{xy}\in\overrightarrow{\text{Aff}(X)},$
where $\overrightarrow{\text{Aff}(X)}$ is the subspace of $\overrightarrow{\mathbb{A}}$ spanned by the vectors $\overrightarrow{xy}$ with $x,y\in X$. So we must have that $\overrightarrow{v}=\overrightarrow{O}$, and (iii) $\Rightarrow$ (i). ∎
Corollary 2.6 has the significant implications below.
Corollary 2.7.
We have that $\mu_{1}=\overrightarrow{O}$ if and only if $\mu_{0}=0$ and either of the equivalent equations
(2.2) $\sum_{z\in X}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})=0,\quad\text{or}\quad\sum_{z\in X\setminus\{x\}}w(z)(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0$
holds for every $x,y\in X$.
Proof.
Equations (2.2) are mere restatements of assertion (iii) in corollary 2.6. The equations are equivalent since
$\sum_{z\in X}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})=\sum_{z\in X\setminus\{x\}}w(z)(\overrightarrow{xz}^{2}-\overrightarrow{yz}^{2})+\left(\sum_{z\in X\setminus\{x\}}w(z)\right)\overrightarrow{xy}^{2}.$
3. Equilibrium Equations
Let $X$ be a finite set of points representing the positions of interacting particles. Let us assume that the force of interaction between two distinct particles $x,y\in X$ can be expressed as
(3.1) $\quad\overrightarrow{F}_{x,y}=\varphi_{x,y}\overrightarrow{xy},\quad\text{with }\ \varphi_{x,y}:\prod_{z\in X}\mathbb{A}\longrightarrow K\ \text{arbitrary}.$
Notice that the expression (3.1) is commonly adopted to describe internal forces in Classical Mechanics, together with conditions on the $\varphi_{x,y}$ in order to comply the homogeneity and isotropy of space [6]. For our purposes, though, it is necessary neither to make these additional assumptions, nor to impose the validity of Newton’s third law of motion, which corresponds to equalities $\varphi_{x,y}=\varphi_{y,x}$ for all $x,y\in X$.
The total force on particle $x$ is $\overrightarrow{F}_{x}=\sum_{y\neq x}\overrightarrow{F}_{x,y}$. As usual, we define an equilibrium of a system of interacting particles as a configuration for which the total force on each particle is zero. For each $x\in X$, let $(X,w_{x})$ be the weighted system defined by
(3.2) $w_{x}(y)=\varphi_{x,y},\quad\text{if}\ y\neq x,\quad w_{x}(x)=-\sum_{y\neq x}\varphi_{x,y}.$
It is clear that each system $(X,w_{x})$ has zero total weight and (constant) first moment $\mu_{1}=\overrightarrow{F}_{x}$. The next result is a direct consequence of the present discussion.
Theorem 3.1.
A configuration $X$ of interacting particles whose forces of interaction are of the form (3.1) is an equilibrium if and only if the corresponding weighted systems $(X,w_{x})$, $x\in X$, defined by equations (3.2) have zero first moment.
Theorem 3.1 represents a landmark in our work as it bridges the theory of moments of weighted systems of section 2 with applications to equilibrium problems. For instance, theorems 3.1, corollary 2.7 and the extended Leibniz identities of appendix B allow us to write equations for equilibria which are homogeneous on the force coefficients $\varphi_{x,y}$ and have a relatively simple form.
Corollary 3.2.
A configuration $X$ is an equilibrium of a systems of particles which interact according to equations (3.1) if and only if, for every pair of points $x,y\in X$, there holds one of the following identities
(3.3) $\sum_{z\in X\setminus\{x\}}\varphi_{x,z}(\overrightarrow{xy}^{2}-\overrightarrow{yz}^{2}+\overrightarrow{zx}^{2})=0,$
or
(3.4) $\sum_{\{z,w\}\subset X}\varphi_{x,z}\varphi_{y,w}\overrightarrow{zw}^{2}=0.$
As a matter of fact, corollary 2.5 ensures that the converse statement in the above corollary is true as long as just the equations (3.4) with $x=y$ hold.
In section 6 we find another set of equilibrium equations for systems which respect Newton’s third law.
The following remarks contain comments on possible applications.
Remark 3.3.
If a function $f$ depends solely on the mutual distances between particles, then the gradient of $f$ for each individual particle has the form (3.1), so corollary 3.2 can be applied to study the critical points of $f$. In Physics, a force field is called conservative when it is the gradient of a potential function. If each $\varphi_{x,y}$ depends only on the distance from $x$ to $y$, that is, if the force fields $F_{x}:\overrightarrow{F}_{x,y}$, $x,y\in X$ are central according to the definition in [6], then $F_{x}$ is shown to be conservative. If, in addition, all the forces derive from a single potential function, a simple calculation shows that Newton’s third law of motion is valid.
Remark 3.4.
External forces (due to a moving reference frame, for instance) acting on the particles may be dealt with by the above method as long as the total force on each particle remains the summation of terms of the form (3.1). The example of central configurations of nonzero masses and nonzero total mass illustrates this possibility, see section 5, equations (5.3) and (5.4) (notice that there, in view of Newton’s second law of motion, we found it more convenient to work with accelerations instead of forces). Reference [7] discusses related problems.
4. The space of weighted systems with $\mu_{1}={\overrightarrow{O}}$
The previous section motivates a deeper study of weighted systems with zero total weight and zero first moment or, more briefly, weighted systems with $\mu_{1}=\overrightarrow{O}$, cf. corollary 2.4. We leave applications and the relation with the second moment respectively to the next section and the appendices.
From now on we let $X=\{x_{1},\ldots,x_{n}\}$. Denote the dimension of the affine closure of $X$ by $\text{dim}(X)$, or just $d$, and call it the dimension of $X$. The difference between the maximum possible value for $\text{dim}(X)$, which is n1n-1, and $d$, that is c=(n1)dc=(n-1)-d, is called the codimension of $X$, and we write $\text{codim}(X)=c$. A weight function $w:X\longrightarrow\mathbb{R}$ may alternatively be represented by the vector $W\in\mathbb{R}^{n}$ whose coordinates are $w(x_{j})=w_{j}$.
Hereafter we shall assume that $X$ is a subset of $\mathbb{R}^{d}$. Select a reference frame at the origin $O\in\mathbb{R}^{d}$, identify points and vectors through $p\leftrightarrow\overrightarrow{Op}$, and use coordinates to represent points and vectors alike. Consider the (augmented) configuration matrix of $X$,
$\mathcal{X}=\left[\begin{array}[]{cccc}1&1&\cdots&1\\ x_{1}&x_{2}&\cdots&x_{n}\end{array}\right].$
Regard $\mathcal{X}$ as a linear map from $\mathbb{W}=\mathbb{R}^{n}$ to $\mathbb{R}^{d+1}$, where $\mathbb{W}$ consists of all possible weight functions (in vector form) defined on $X$. We call $\mathbb{W}$ the weight space of $X$. The kernel of $\mathcal{X}$ is the subspace consisting of weight vectors such that $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{O}$, that is, the subspace
$\mathbb{W}_{0}(X)=\{W\in\mathbb{W}\,\big/\,\ \mu_{1}=\overrightarrow{O}\}.$
The main objective in this section is to describe $\mathbb{W}_{0}(X)$.
Lemma 4.1.
$\dim\mathbb{W}_{0}(X)=\text{\emph{codim}}\,(X)$.
Proof.
The configuration matrix $\mathcal{X}$ has the same rank as the matrix
$\left[\begin{array}[]{cccc}1&0&\cdots&0\\ x_{1}&x_{2}-x_{1}&\cdots&x_{n}-x_{1}\end{array}\right],$
that is, $d+1$. The nullity of $\mathcal{X}$ is thus n(d+1)=(n1)d=cn-(d+1)=(n-1)-d=c. This is the dimension of $\mathbb{W}_{0}(X)$. ∎
Theorem 4.2.
$\bigwedge^{\text{\emph{codim}}\,(X)+1}\mathbb{W}_{0}(X)=\{O\}$.
Proof.
Recall $\bigwedge$ represents the exterior product of vector spaces. If a vector space has finite dimension, then its $k$-th exterior power is equal to the trivial space for all $k$ greater than the dimension of the vector space [20]. ∎
Many identities involving the values of distinct weight functions arise from theorem 4.2. This topic will be further discussed in section 5.
Next we present a systematic method for constructing bases for $\mathbb{W}_{0}(X)$.
As we shall soon see, a key role is played by configurations whose codimension is 1. Borrowing the terminology from Celestial Mechanics, such a configuration will be designated as Dziobek configuration, or Dziobek subconfiguration, when it is a proper subset of a larger configuration.
Notice that a configuration $X$ has dimension $d$ if and only if $X$ contains an ascending chain of simplexes $S_{0}\subset\ldots\subset S_{d-1}\subset S_{d}$, where each $S_{k}$ is $k$-dimensional, $k=0,\ldots,d-1,d$. We refer to any simplex of dimension $d$ contained in $X$ as a core of $X$. There are n(d+1)=cn-(d+1)=c points in $X\setminus S_{d}$, so we have exactly $c$ distinct Dziobek subconfigurations $D^{(1)},\ldots,D^{(c)}$ containing the core $S_{d}$. Figure 1 depicts one such arrangement of simplexes and Dziobek subconfigurations of $X$. We call each such diagram a Dziobek tree for $X$.
Consider a Dziobek tree for $X$. If $c=0$, we have that $X$ is the core $S_{d}$ and, by proposition 4.1, $\mathbb{W}_{0}(X)$ is the trivial space. The cases $c>0$ are dealt with in the next proposition. For each Dziobek subconfiguration $D^{(l)}$ containing $S_{d}$, let us embed the corresponding $\mathbb{W}_{0}(D^{(l)})$, $l=1,\ldots,c,$ as a (one-dimensional) subspace of $\mathbb{W}_{0}(X)$ through the map
(4.1) $\iota_{l}:\mathbb{W}_{0}(D^{(l)})\hookrightarrow\mathbb{W}_{0}(X),$
which is defined by placing zeros at the coordinates corresponding to the n(d+2)=c1n-(d+2)=c-1 points in $X\setminus D^{(l)}$.
Figure 1. Sketch of a Dziobek tree. The “trunk” is the ascending chain of $k$-dimensional simplexes $S_{k}$, $k=0,\ldots,d$, and the “canopy” consists of all Dziobek subconfigurations $D^{(l)}$, $l=1,\ldots,c$, containing the core $S_{d}$.
Proposition 4.3.
To every core $S_{d}$ of a configuration $X$ of codimension $c>0$ there corresponds a unique direct sum decomposition
$\mathbb{W}_{0}(X)=\bigoplus_{l=1}^{c}\iota_{l}\left(\mathbb{W}_{0}(D^{(l)})\right),$
where $D^{(l)}$, $l=1,\ldots,c$, are the Dziobek subconfigurations of $X$ which contain $S_{d}$.
Proof.
After possibly reindexing the elements of $X$, we may assume that the core is $S_{d}=\{x_{1},\ldots,x_{d+1}\}$. Let $D^{(l)}=S_{d}\cup\{x_{d+l+1}\}$, $l=1,\ldots,c$. $D^{(l)}$ has codimension 1, so $\mathbb{W}_{0}(D^{(l)})$ is spanned by a weight vector $(w_{1},\ldots,w_{d+1},w_{d+l+1})$ such that
(4.2) $\sum_{j=1}^{d+1}w_{j}+w_{d+l+1}=0,\qquad\sum_{j=1}^{d+1}w_{j}x_{j}+w_{d+l+1}x_{d+l+1}=O.$
We must have that $w_{d+l+1}\neq 0$, for all $l=1,\ldots,c$, otherwise the above equations would provide a nontrivial linear combination of the vectors $\overrightarrow{x_{1}x_{j}}$, $j=2,\ldots,d+1$, contradicting the hypothesis that $S_{d}$ is a core. By adding c1c-1 zeros as previously explained, we embed each of the spanning vectors for the Dziobek subconfigurations $D^{(l)}$ into $\mathbb{W}_{0}(X)$, thus obtaining the collection of $c$ weight vectors
$(*,w_{d+2},0,\ldots,0),\ (*,0,w_{d+3},\ldots,0),\ \ldots,\ (*,0,\ldots,0,w_{n}),$
where each * represents a (nonzero) vector in $\mathbb{R}^{d+1}$. This collection is clearly a basis for $\mathbb{W}_{0}(X)$. ∎
Remark 4.4.
Equations (4.2) characterize $w_{1},\ldots,w_{d+1}$ as the homogeneous barycentric coordinates of $x_{d+l+1}$ with respect to the simplex $S_{d}$, that is to say, $x_{d+l+1}=\text{bar}\left(S_{d},w\big|_{S_{d}}\right)$. Indeed, equations (4.2) imply that
$(w_{1}+\ldots+w_{d+1})x_{d+l+1}=w_{1}x_{1}+\ldots w_{d+1}x_{d+1},$
with $w_{1}+\ldots+w_{d+1}\neq 0$. Compare with part (2) of proposition 2.3.
Considering all possible Dziobek trees for $X$, we see from proposition 4.3 that the maximum number of bases obtainable by our method is ${n\choose d+1}={n\choose c}$. The maximum is attained when each subconfiguration of $X$ formed by $d+1$ points is a core of $X$. In this case, we call $X$ a generic configuration. A configuration is termed nongeneric if it is not generic. The type of a nongeneric configuration refers to the number of its subconfigurations with $d+1$ points that are not cores. We illustrate the possibilities for five points on a plane in figure 2. An application is made in example 5.3.
Figure 2. Planar configurations of five points: generic (left), nongeneric of type 11 (left center), nongeneric of type 2 (center right), and nongeneric of type 44 (right).
For the case $c=1$, a generator with geometric meaning for $\mathbb{W}_{0}(X)$ can be obtained using basic properties of determinants. The following proposition is proved in [28].
Proposition 4.5.
Let $X=\{x_{1},\ldots,x_{n}\}$, and let $X_{j}$ be the subconfiguration $X\setminus\{x_{j}\}$. Denote by $\mathcal{X}_{j}$ the configuration matrix of $_{j}. XX$ has codimension $1$ if and only if $\mathbb{W}_{0}(X)$ is spanned by the weight vector
$W=(-\Delta_{1},\Delta_{2},-\Delta_{3},\ldots,(-1)^{n}\Delta_{n}),$
where $\Delta_{j}=\det(\mathcal{X}_{j})$.
The cases $c>1$ require a definition and additional notation.
The (oriented) volume of a configuration $X$ of codimension $c$ is the determinant $\det(\mathcal{X})$ of its configuration matrix $\mathcal{X}$ if $c=0$, and zero otherwise. Let $X_{j_{1}\ldots j_{\beta}}$ denote the subconfiguration $X\setminus\{x_{j_{1}},\ldots,x_{j_{\beta}}\}$, where $j_{1}<\ldots<j_{\beta}$, and let $\mathcal{X}_{j_{1}\ldots j_{\beta}}$ be the configuration matrix of $X_{j_{1}\ldots j_{\beta}}$. Denote the volume of $X_{j_{1}\ldots j_{\beta}}$ by $\Delta_{j_{1}\ldots j_{\beta}}$. The construction of bases for $\mathbb{W}_{0}(X)$ using the volume barycentric coordinates $\Delta_{j_{1}\ldots j_{\beta}}$ is similar to the one done in the proof of proposition 4.3 and employs the embeddings $\iota_{l}$, see (4.1). For the sake of concreteness, we provide an example.
Example 4.6.
Consider seven points $x_{1},\ldots,x_{7}$ forming a two-dimensional configuration $X$. So $X$ has codimension $c=4$. Let us determine a basis for $\mathbb{W}_{0}(X)$ using volumes and proposition 4.5.
Suppose $X_{4567}=\{x_{1},x_{2},x_{3}\}$ is a core of $X$; the corresponding volume $\Delta_{4567}$ is, thus, nonzero. $X_{4567}$ is contained in four Dziobek subconfigurations: $X_{567}$, $X_{467}$, $X_{457}$, $X_{456}$. These Dziobek subconfigurations produce the respective weight vectors
$W_{4}$ $=(-\Delta_{1567},\Delta_{2567},-\Delta_{3567},\Delta_{4567},0,0,0),$
$W_{5}$ $=(-\Delta_{1467},0,\Delta_{2467},-\Delta_{3467},0,\Delta_{4567},0,0),$
$W_{6}$ $=(-\Delta_{1457},\Delta_{2457},0,-\Delta_{3457},0,0,\Delta_{4567},0),$
$W_{7}$ $=(-\Delta_{1456},\Delta_{2456},-\Delta_{3456},0,0,0,\Delta_{4567}),$
where the index of each vector is the index of the point added to $X_{4567}$ in order to form the corresponding Dziobek subconfiguration. Since $\Delta_{4567}\neq 0$, clearly $W_{4},W_{5},W_{6},W_{7}$ form a basis of $\mathbb{W}_{0}(X)$.
Now, if $X_{4567}$ is not a core of $X$, then it necessarily has dimension one, and we can replace one of its points with one of the points in $X\setminus X_{4567}$, say replace $x_{3}$ with $x_{4}$. If $X_{3567}$ is a core, we construct a basis using our procedure, otherwise, we make another replacement. Such a sequence of replacements will eventually produce a core of $X$, or else $X$ is collinear, which is a contradiction. Notice that the same argument shows that replacements can be used to successively obtain each element $S_{k}$ of a Dziobek tree for $X$.
We end this section with a criterion for membership in $\mathbb{W}_{0}(X)$. The proof is followed by related remarks.
Theorem 4.7.
Let $X$ be a configuration of dimension $d$ such that $\Delta_{d+2\ldots n}\neq 0$. We have that $(w_{1},\ldots,w_{n})\in\mathbb{W}_{0}(X)$ if and only if the identities
(4.3) $(-1)^{i-d}w_{i}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}w_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\quad i=1,\ldots,d+1,$
hold, where $\widehat{l}$ indicates the removal of the index $l$.
Proof.
By hypothesis, the subconfiguration $\{x_{1},\ldots,x_{d+1}\}$ is a core of $X$. Thus we have the following basis of $\mathbb{W}_{0}(X)$:
$(-\Delta_{1d+3\,d+4\ldots n},\ldots,(-1)^{d+1}\Delta_{d+1\,d+3\,d+4\ldots n},(-1)^{d+2}\Delta_{d+2\,d+3\,d+4\ldots n},0,\ldots,0),$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \$ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ $\vdots$
$(-\Delta_{1d+2\,d+3\ldots n-1},\ldots,(-1)^{d+1}\Delta_{d+1\,d+2\,d+3\ldots n-1},0,\ldots,0,(-1)^{d+2}\Delta_{d+2\,d+3\,d+4\ldots n}),$
Each of the above weight vectors corresponds to a Dziobek subconfiguration $\{x_{1},\ldots,x_{d+1},$ $x_{l}\}$, with $l=d+2,\ldots,n$ ; we denote the weight vector associated to $l$ by $W_{l}$.
If $W=(w_{1},\ldots,\ldots,w_{n})\in\mathbb{W}_{0}(X),$ we have that $W=\sum_{d+2}^{n}a_{l}W_{l}$, with
$a_{l}=\frac{w_{l}}{(-1)^{d+2}\Delta_{d+2\ldots n}},\quad l=d+2,\ldots,n.$
On the other hand, for $i=1,\ldots,d+1$, we have that
(4.4) $w_{i}=\sum_{l=d+2}^{n}a_{l}(-1)^{i}\Delta_{i\,d+2\ldots\widehat{l}\ldots n}=\sum_{l=d+2}^{n}(-1)^{i-d}w_{l}\frac{\Delta_{i\,d+2\ldots\widehat{l}\ldots n}}{\Delta_{d+2\ldots n}},$
and the second equality promptly implies equations (4.3).
For the converse, given a weight vector $(w_{1},\ldots,w_{n})$, if we choose $a_{l}$ so that
$w_{l}=(-1)^{d+2}a_{l}\Delta_{d+2\ldots n},\ \ \ l=d+2,\ldots,n,$
then equations (4.3) imply that
$\Delta_{d+2\ldots n}\left(w_{i}-\sum_{l=d+2}^{n}(-1)^{i}a_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n}\right)=0,\ \ \ i=1,\ldots,d+1.$
Since $\Delta_{d+2\ldots n}\neq 0$, we obtain
$w_{i}=\sum_{l=d+2}^{n}(-1)^{i}a_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\ \ \ i=1,\ldots,d+1.$
Hence $(w_{1},\ldots,w_{n})=\sum_{d+2}^{n}a_{l}W_{l}$, that is, $W\in\mathbb{W}_{0}(X).$
Remark 4.8.
The first implication in theorem 4.7 does not require the hypothesis $\Delta_{d+2\ldots n}\neq 0$. We illustrate the argument, which can be found in [14, 28], in the context of central configurations.
If $(w_{1},\ldots,w_{n})\in\mathbb{W}_{0}(X)$, then
$\sum_{j\neq k}w_{j}\overrightarrow{x_{k}x_{j}}=\overrightarrow{O},\quad k=1,\ldots,n.$
Set $k=1$ and take the exterior product with $\overrightarrow{x_{1}x_{3}}\wedge\ldots\wedge\overrightarrow{x_{1}x_{d+1}}$. Let $\overrightarrow{e}_{1},\ldots,\overrightarrow{e}_{d}$ be the canonical basis of $\mathbb{R}^{d}$. From
$\Delta_{d+2\ldots n}\,\overrightarrow{e}_{1}\wedge\ldots\overrightarrow{e}_{d}$ $=\overrightarrow{x_{1}x_{2}}\wedge\overrightarrow{x_{1}x_{3}}\wedge\ldots\wedge\overrightarrow{x_{1}x_{d+1}},\ \ \text{and}$
$\Delta_{2\,d+2\ldots\widehat{l}\ldots n}\,\overrightarrow{e}_{1}\wedge\ldots\overrightarrow{e}_{d}$ $=(-1)^{1-d}\overrightarrow{x_{1}x_{l}}\wedge\overrightarrow{x_{1}x_{3}}\wedge\ldots\wedge\overrightarrow{x_{1}x_{d+1}},\ \ \text{if }\ l>d+1,$
we get
$(-1)^{2-d}w_{2}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}w_{l}\Delta_{2\,d+2\ldots\widehat{l}\ldots n},$
which is (4.3) for $i=2$.
Remark 4.9.
The choice of the core used in theorem 4.7 was just a matter of convenience. We could have chosen any other core of $X$ and determined formulæ  corresponding to (4.3) by following the same procedure. We also notice that equations (4.3) can be obtained using theorem 4.2, albeit not in as straightforward a way.
Remark 4.10.
Consider a Dziobek subconfiguration which does not contain the core used in theorem 4.7. Interestingly, if we write equations (4.3) for the weight vector associated to this Dziobek subconfiguration, we obtain identities corresponding to the Plücker relations used in Algebraic Geometry to describe Grassmannians as projective varieties [24, 25]. For instance, if $n=5$ and $d=2$, consider the two-dimensional subspace spanned by the rows of the $2\times 4$ matrix whose columns are $\overrightarrow{x_{1}x_{2}},\ldots,\overrightarrow{x_{1}x_{5}}$. The Plücker coordinates of this subspace are $p_{12}=\Delta_{23}$, $p_{13}=\Delta_{35}$, $p_{14}=\Delta_{34}$, $p_{23}=\Delta_{25}$, $p_{24}=\Delta_{24}$ and $p_{34}=\Delta_{45}$. If $X$ has core $\{x_{1},x_{2},x_{3}\}$, then $\Delta_{45}\neq 0$, and we can consider the basis of $\mathbb{W}_{0}(X)$ formed by
$W_{4}=(-\Delta_{14},\Delta_{24},-\Delta_{34},0,\Delta_{45}),\quad W_{5}=(-\Delta_{15},\Delta_{25},-\Delta_{35},\Delta_{45},0).$
If we take the weight vector associated to the Dziobek subconfiguration $\{x_{1},x_{2},x_{4},$ $x_{5}\}$, which is $(-\Delta_{13},\Delta_{23},0,-\Delta_{34},$ $\Delta_{35})$, and apply formula (4.3) with $i=2$, we find that
$\Delta_{23}\Delta_{45}-\Delta_{24}\Delta_{35}+\Delta_{25}\Delta_{34}=0,$
which corresponds to the well known unique Plücker relation $p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}=0$ defining the Grassmannian of two-dimensional subspaces of $\mathbb{R}^{4}$ as a variety in $\mathbb{R}P^{5}$.
5. Some Applications to central configurations
We call a configuration of particles (point masses, bodies) in $\mathbb{R}^{N}$ any weighted system $(X,m)$ where $m$ is a nonvanishing function which we traditionally refer to as mass. Recall the notation used in sections 3 and 4. The points of $X$ are denoted by $x_{i}$, $m_{i}$ is the value of $m$ at $x_{i}$, $i=1,\ldots,n$, the dimension of $X$ is $d$ and the codimension of $X$ is c=(n1)dc=(n-1)-d. We may interpret $m$ as the weight vector $(m_{1},\ldots,m_{n})\in\mathbb{W}=\mathbb{R}^{n}$. Assuming the particles interact under some homogeneous potential, the corresponding acceleration vectors have the form
$\overrightarrow{\gamma}_{j}=-\sum_{i\neq j}m_{i}s_{ij}^{a}\overrightarrow{x_{i}x_{j}},\quad s_{ij}=\overrightarrow{x_{i}x_{j}}^{2},\ \ a\in\mathbb{R},$
according to Newton’s second law of motion.
Following Albouy [4], we say that $(X,m)$ is a central configuration if there exist a vector $\overrightarrow{\gamma}_{O}$, a point $x_{O}$ and a real number $\lambda$ such that
(5.1) $\overrightarrow{\gamma}_{j}-\overrightarrow{\gamma}_{O}=\lambda\overrightarrow{x_{O}x_{j}},\quad j=1,\ldots,n.$
Albouy has shown that, if $a=0$, then any configuration is central, and if $a\neq 0$ and d=n1d=n-1, i.e, if $c=0$, then $X$ is a regular simplex for all values of $\mu_{0}$. In the cases $\mu_{0}\neq 0$, this assertion follows from lemma 4.1 and theorem 5.2 below (as in [4], we tacitly assume that $\lambda/\mu_{0}>0$).
From now on we suppose that $a\neq 0$.
Multiplying both sides of (5.1) by $m_{j}$ and summing over $j$, we obtain an equation relating the constants $\lambda,x_{O},\overrightarrow{\gamma}_{O}$ and the total mass $\mu_{0}$, namely
(5.2) $\lambda\mu_{1}(x_{O})=-\mu_{0}\overrightarrow{\gamma}_{O}.$
Proposition 5.1.
Let $(X,m)$ be a central configuration with $\mu_{0}=0$ and $\lambda\neq 0$. Let $X_{j}=X\setminus\{x_{j}\}$, $j=1,\ldots,n$. We have that
  • (1)
    $m\in\mathbb{W}_{0}(X)$.
  • (2)
    $x_{j}=\text{\emph{bar}}\left(X_{j},m\big|_{X_{j}}\right)$, for every $j=1,\ldots,n$.
  • (3)
    $X$ has codimension $c\geq 1$ and $\mu_{2}$ is constant. In particular, we have the Leibniz identity $\sum_{i<j}m_{i}m_{j}s_{ij}=0$, and
    $\sum_{k=1}^{n}m_{k}(s_{ij}-s_{jk}+s_{ik})=0,\quad\text{for every }\ i,j=1,\ldots,n.$
    If $c=1$ then, for every $j=1,\ldots,n$, $X_{j}$ is a $d$-dimensional simplex (i.e., $X_{j}$ is a core of $X$) and $x_{j}$ is not on the boundary of the convex hull of $X_{j}$ (so $X$ is a generic configuration).
  • (4)
    $X$ has dimension $d\leq n-2$, and if $\Delta_{d+2\ldots n}\neq 0$, then
    $(-1)^{i-d}m_{i}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}m_{l}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\quad i=1,\ldots,d+1,$
Proof.
Part (1) follows from equation (5.2). (2) is a direct consequence of proposition 2.3, because $m$ is nonvanishing. In order to prove (3), note that $\mu_{1}=\overrightarrow{O}$ provides a nontrivial linear relation for the elements of $X$. Hence $c\geq 1$. Moreover, $\mu_{2}$ is constant according to corollary 2.6. The equations follow from corollaries 2.5 and 2.7. The assertion on the case $c=1$ follows directly from (1), (2), and the barycentric interpretation of the (nonzero) coordinates of the mass vector. Compare with proposition 4.5. Finally, (4) is immediate from part (1) and theorem 4.7. ∎
In the case $n=4$, a full classification of a special family of central configurations with $\mu_{0}=0$, including examples of central configurations with $\lambda\neq 0$, can be found in [11].
Next, let us consider the case $\mu_{0}\neq 0$. As proved in [4], $(X,m)$ is a central configuration if and only if $\overrightarrow{\gamma}_{O}$ is the null vector and $x_{O}=\text{bar}(X,m)$ is the center of mass of $(X,m)$. The central configuration equations can thus be put in the form
(5.3) $\sum_{i\neq j}m_{i}S_{ij}\overrightarrow{x_{j}x_{i}}=\overrightarrow{O},\quad j=1,\ldots,n,$
where $S_{ij}=s_{ij}^{a}-\frac{\lambda}{\mu_{0}}$. Equations (5.3) imply that, for each $j$, a weighted system $(X,C_{j})$ with $\mu_{1}=\overrightarrow{O}$ can be defined, namely
(5.4) $C_{j}(x_{i})=m_{i}S_{ij},\quad j\neq i,\ \ C_{j}(x_{j})=-\sum_{i\neq j}m_{i}S_{ij},\ \ \text{for}\ \ j=1,\ldots,n.$
For convenience, we write $C_{j}(x_{j})=m_{j}S_{jj}$. We have just proved the next theorem.
Theorem 5.2.
A configuration of particles $(X,m)$ with $\mu_{0}\neq 0$ is a central configuration if and only if each weighted systems $(X,C_{j})$, with $C_{j}$ defined by equations (5.4), $j=1,\ldots,n$, have zero first moment.
Therefore we can apply the theory developed in sections 2 and 4 to the study of central configurations with $\mu_{0}\neq 0$. In particular, we recover and extend some known results. Additional results on central configurations are provided in section 6 and appendix B.
Corollary (Albouy-Chenciner).
A configuration of particles $(X,m)$ with $\mu_{0}\neq 0$ is central if and only if, for each $j,k=1,\ldots,n$, there holds
$\sum_{i\neq j}m_{i}S_{ij}(s_{ij}+s_{jk}-s_{ki})=0.$
The above formulæ  were found by the authors of the celebrated reference [3] and have since been named after them. Reference [22] contains an alternative deduction.
The equations in the next corollary correspond to the formulæ  in theorem 2.2 of [14].
Corollary (Dias).
$(X,m)$ is a central configuration with $\mu_{0}\neq 0$, of dimension $d\leq n-2$ and such that $\Delta_{d+2\ldots n}\neq 0$ if and only if
$(-1)^{i-d}m_{i}S_{ij}\Delta_{d+2\ldots n}=\sum_{l=d+2}^{n}m_{l}S_{lj}\Delta_{i\,d+2\ldots\widehat{l}\ldots n},\quad i=1,\ldots,d+1,\ j=1,\ldots,n.$
Corollary (Extended Dziobek).
Let $(X,m)$ be a central configuration with $\mu_{0}\neq 0$ and codimension $c$. Every $(c+1)\times(c+1)$-minor of any $n\times(k+l)$-matrix, with $k+l=c+1$, formed by $k$ central configuration weight vectors $C_{j}$ (see equations (5.4)) and any $l$ weight vectors in $\mathbb{W}_{0}(X)$ (for instance, those corresponding to Dziobek subconfigurations of $X$, see section 4) vanishes. In particular, any $(c+1)\times(c+1)$-minor of the $n\times n$ matrix formed by the $S_{ij}$ is zero.
The previous corollary follows immediately from theorems 4.25.2 and the fact that $m$ is nonvanishing. In the Dziobek case, $c=1$, the last statement in the corollary implies that $S_{ij}S_{kl}-S_{il}S_{jk}=0$ for any four indices $i,j,k,l$. When the indices are distinct, these identities are the well known equations (18) in [28]. By considering some equal indices in the case $n=4$ and $c=1$, for instance, we compute $S_{12}S_{24}-S_{14}S_{22}=0$, which gives
$m_{1}S_{12}S_{14}+m_{2}S_{12}S_{24}+m_{3}S_{14}S_{23}+m_{4}S_{14}S_{24}=0.$
Thus, if all the masses have the same sign, then $S_{12},S_{14},S_{23}$ and $S_{24}$ cannot have the same sign. For central configurations with $c=2$, e.g. those formed by five bodies on a plane or six bodies in space, we have that
$\left|\begin{array}[]{ccc}S_{i_{1}j_{1}}&S_{i_{1}j_{2}}&S_{i_{1}j_{3}}\\ S_{i_{2}j_{1}}&S_{i_{2}j_{2}}&S_{i_{2}j_{3}}\\ S_{i_{3}j_{1}}&S_{i_{3}j_{2}}&S_{i_{3}j_{3}}\end{array}\right|=0,$
for any choice of indices.
The next example illustrates the use of determinants provided by theorem 4.2 to obtain equations tailored for particular families of central configurations.
Example 5.3.
Consider a planar configuration $X=\{x_{1},x_{2},x_{3},x_{4},x_{5}\}$ and let $x_{i}x_{j}$ denote the line through the points $x_{i}$ and $x_{j}$. Suppose $x_{3}\in x_{1}x_{2}$ and $x_{4},x_{5}\notin x_{1}x_{2}$. In the terminology of section 4, $X$ is a nongeneric configuration of type 1 (see figure 2, page 2).
We have five weight vectors associated to the Dziobek subconfigurations $X_{1},\ldots,$ $X_{5}$ of $X$, namely
$W_{1}$ $=(0,-\Delta_{12},\Delta_{13},-\Delta_{14},\Delta_{15}),\quad W_{2}=(-\Delta_{12},0,\Delta_{23},-\Delta_{24},\Delta_{25})$
$W_{3}$ $=(-\Delta_{13},\Delta_{23},0,-\Delta_{34},\Delta_{35}),\quad W_{l}=(-\Delta_{1l},\Delta_{2l},-\Delta_{3l},0,0),\ \ l=4,5.$
Theorems 4.2 and 5.2 imply that, if $i,j,k=1,2,3$ and $l=4,5$, then
(5.5) $\left|\begin{array}[]{ccc}m_{3}S_{3j}&*&-\Delta_{3l}\\ m_{4}S_{4j}&-\Delta_{i4}&0\\ m_{5}S_{5j}&\Delta_{i5}&0\end{array}\right|=0\ \ \text{and}\ \ \left|\begin{array}[]{ccc}m_{i}S_{ij}&*&(-1)^{i}\Delta_{il}\\ m_{k}S_{kj}&*&(-1)^{k}\Delta_{kl}\\ m_{4}S_{4j}&-\Delta_{i4}&0\end{array}\right|=0,$
for any central configuration $(X,m)$ with $\mu_{0}\neq 0$. We have the following sets of necessary conditions for such $(X,m)$.
Proposition 5.4.
Suppose $(X,m)$ is a planar central configuration of five particles with nonzero total mass $\mu_{0}$, and such that $x_{3}\in x_{1}x_{2}$ and $x_{4},x_{5}\notin x_{1}x_{2}$. Let $h_{l}$ be the distance from $x_{l}$ to $x_{1}x_{2}$.
  • (a)
    For every $j=1,2,3,$ we have that
    $m_{4}S_{j4}h_{4}=\pm m_{5}S_{j5}h_{5},$
    where the ++ (resp. -) sign occurs if $x_{4}$ and $x_{5}$ are on opposite sides (resp. the same side) of $x_{1}x_{2}$.
  • (b)
    Suppose $x_{4}$ and $x_{5}$ are on opposite sides of $x_{1}x_{2}$. If $h_{4}=h_{5}$, and all the masses have the same sign, then $x_{1}x_{2}$ is the perpendicular bisector of $x_{4}x_{5}$, and $m_{4}=m_{5}$.
  • (c)
    Suppose $x_{4}$ and $x_{5}$ are on the same side of $x_{1}x_{2}$. If $s_{24}^{a}$ or $s_{25}^{a}$ equal $\lambda/\mu_{0}$, then $s_{24}^{a}=s_{25}^{a}=\lambda/\mu_{0}$, and
    $m_{1}S_{12}s_{12}^{1/2}=m_{3}S_{23}s_{23}^{1/2}.$
    In particular, for a<1/2a<-1/2 and $m_{1}=m_{3}$, particle $x_{2}$ is at the midpoint of $x_{1}$ and $x_{3}$. Moreover, if the lines $x_{1}x_{2}$ and $x_{4}x_{5}$ are parallel, then, for all $m_{2}$, the configuration is an isosceles trapezoid and $m_{4}=m_{5}$.
Proof.
Item (a) follows from the first determinant in equations (5.5) and from $\frac{\Delta_{j4}}{\Delta_{j5}}=\pm\frac{h_{5}}{h_{4}}$. For (b), notice $h_{4}=h_{5}$ forces $x_{1}x_{2}$ to intersect $x_{4}x_{5}$ at the midpoint of $x_{4}$ and $x_{5}$. Apply the Perpendicular Bisector Theorem [4] and item (a) to draw the stated conclusions. To obtain the equation in (c), we substitute $S_{24}=0$ in the second determinant of (5.5) with $i=1,j=2,k=3,l=4$ and use the equation $\frac{\Delta_{14}}{\Delta_{34}}=\frac{s_{23}^{1/2}}{s_{12}^{1/2}}$. The remaining statements in (c) follow from the fact that $f(x)=x^{2a+1}-(\lambda/\mu_{0})x$, $x>0$, is an injective function for any a<1/2a<-1/2, and from item (a). Indeed, $h_{4}=h_{5}$ by hypothesis, so $X$ is an isosceles trapezoid and, by symmetry, $S_{14}=S_{35}$ and $S_{34}=S_{15}$. Thus, by (a),
$m_{5}S_{15}=-m_{4}S_{14}=-m_{4}S_{35}=m_{4}\left(-\frac{m_{4}}{m_{5}}S_{34}\right)\ \Longrightarrow\ m_{4}=\pm m_{5}.$
But $m_{4}=-m_{5}$ implies $S_{14}=S_{15}$, which is not possible. ∎
References [12, 18] describe a family of planar central configurations of five bodies under the Newtonian gravitational potential (i.e., for a=3/2a=-3/2) which fulfills the conditions of item (c) of proposition 5.4. In both [12] and [18], the masses are positive and the symmetry of $(X,m)$ is assumed from start. On the other hand, the hypothesis $s_{24}^{a}=\lambda/\mu_{0}$ is part of the conclusions in [12], which also discusses the changes in the shape of the isosceles trapezoid as $m_{2}$ changes. We note that reference [18] examines configurations considered in item (b) of proposition 5.4, as well as some symmetrical configurations with $m_{4}\neq m_{5}$ (none of which turns out to be central).
6. Generalized Dziobek Equations
Let $\mathfrak{F}=\{(X,f_{j})\}_{j=1}^{n}$, be a family of weighted systems with configuration $X$. We define the weight matrix of $\mathfrak{F}$ as the $n\times n$ matrix $\mathcal{F}$ whose entries are $\mathcal{F}_{ij}=f_{j}(x_{i})$, $i,j=1,\ldots,n$. Suppose each $(X,f_{j})\in\mathfrak{F}$ has $\mu_{1}=\overrightarrow{O}$. If $W_{1},\ldots,W_{c}$ is a basis of $\mathbb{W}_{0}(X)$, there exists a $c\times n$ matrix $\mathcal{A}$ such that
(6.1) $\mathcal{F}=\mathcal{W}\mathcal{A},$
where $\mathcal{W}$ is the matrix whose columns are the basis vectors.
When $\mathcal{F}$ is a symmetric matrix, we refer to the family $\mathfrak{F}$ as a symmetric family of weighted systems. Examples of such families are the systems of interacting particles that obey Newton’s third law of motion, see section 3. Clearly
$\mathcal{F}=\mathcal{F}^{T}\quad\text{if and only if}\quad\mathcal{W}\mathcal{A}=\mathcal{A}^{T}\mathcal{W}^{T}.$
Let $\mathcal{S}$ be the $c\times c$ matrix such that $\mathcal{A}=\mathcal{S}\mathcal{W}^{T}$; notice that $\mathfrak{F}$ is a symmetric family if and only if $\mathcal{S}$ is symmetric. So, if $\mathfrak{F}$ is symmetric, we can rewrite equation (6.1) as
(6.2) $\mathcal{F}=\mathcal{W}\mathcal{S}\mathcal{W}^{T}.$
Conversely, if we set $\mathcal{S}\mathcal{W}^{T}=\mathcal{A}$ in equation (6.2), we recover equation (6.1). Thus the following statement has been proved.
Theorem 6.1.
The weight vectors of a symmetric family $\mathfrak{F}=\{(X,f_{j})\}_{j=1}^{n}$ of weighted systems with a configuration $X=\{x_{1},\ldots,x_{n}\}$ are in $\mathbb{W}_{0}(X)$ if and only if equation (6.2) is satisfied by the weight matrix $\mathcal{F}=(f_{j}(x_{i}))_{1\leq i,j\leq n}$ of $\mathfrak{F}$, for some basis of $\mathbb{W}_{0}(X)$ and some symmetric matrix $S$.
Remark 6.2.
Equation (6.2) contains the classical Dziobek equations for central configurations of codimension one [4, 16, 29], so we call (6.2) the generalized Dziobek equations. Observe that, in order to apply theorem 6.1 to central configurations, we must take $\mathcal{F}=\mathcal{C}\mathcal{M}$, where $\mathcal{C}$ is the matrix whose $j^{\text{th}}$ column is the weight vector $C_{j}$ defined by (5.4), and $\mathcal{M}$ is the diagonal matrix with entries $m_{1},\ldots,m_{n}$ on the main diagonal. Remark 3.3 indicates other possible applications of theorem 6.1.
Remark 6.3.
The number of parameters defining $S$ is $c+1\choose 2$}, which, according to formula (A.2.4) in appendix A, is precisely the number of constraints that must be imposed on the mutual distances of a configuration of $n$ points to guarantee it has a given codimension $c>0$. This coincidence is not fortuitous since, by theorem A.2.2, the vanishing of $c+1\choose 2$} Cayley-Menger determinants provide the necessary constraints. Dziobek’s approach via critical point theory then introduces $c+1\choose 2$} Lagrange multipliers. Dziobek also verified that the derivative of a Cayley-Menger determinant with respect to a squared mutual distance can be expressed simply as a constant times the product of a pair of oriented volumes of suitable subconfigurations [16, 21]. Aside from determining the exact value of the multiplicative constant, this fact can be readily inferred from theorem 6.1. Our proof of theorem 6.1 was inspired by the elegant proof of proposition 8 in [4].
Appendix A Constraints for Mutual Distances
The HLK theorem and corollary 2.6 play a relevant role in obtaining constraints for mutual distances of configurations. We separate these constraints into two sets: the direct (and simpler) ones that follow straight from the HLK theorem, and the constraints obtained by taking determinants of special matrices associated with a given configuration.
A.1. Direct constraints
Early on in the investigations towards this manuscript, we came across the following geometric fact: the squared distances $s_{i}$, $i=1,\ldots,4$ from the vertices of a rhombus of squared semidiagonals $s=\overrightarrow{x_{1}x_{3}}^{2}$ and $s^{\prime}=\overrightarrow{x_{2}x_{4}}^{2}$ to any point in space satisfy the constraint
$s_{1}-s_{2}+s_{3}-s_{4}=2(s-s^{\prime}).$
Notice the expression on the right-hand side of the above equation is the value of the left-hand side at the centroid of the rhombus. This observation led us to search for simpler alternatives to the current methods for obtaining constraints for mutual distances of configurations, specially to the Cayley-Menger determinant conditions soon to be discussed. We encountered the following corollary to the HLK theorem.
Corollary A.1.1.
Let $(X,w)$ and $(X^{\prime},w^{\prime})$ be weighted systems in an Euclidean affine space whose total weights are nonzero and equal, and whose barycenters coincide. The difference between the respective second moments has the same value at all points, namely, the value it has at the (common) barycenter.
Proof.
We just have to apply part (1) of the HLK theorem to $(X,w)$ and $(X^{\prime},w^{\prime})$, subtract the corresponding sides of the resulting equations and do a simple manipulation. ∎
Corollary A.1.1 allows us to generalize our earlier remarks on the rhombus in many ways. For instance, it implies that the difference between the sums of the squared distances from any point to the vertices of two concentric regular $n$-gons of respective squared radii $s$ and $s^{\prime}$ is always equal to $n(s-s^{\prime})$, that is, it depends neither on the point nor on the relative position of the $n$-gons.
A.2. Determinantal constraints
In the classical paper [10], Cayley established determinantal conditions for the dimension of a configuration and for its co-sphericity. Cayley’s conditions will be easily deduced using weighted systems and their moments. We shall also provide a theorem which compares the kernels of the matrices involved. Finally, we prove a theorem which exhibits a complete set of independent constraints on mutual distances to ensure that a configuration has a given dimension.
Given a configuration $X=\{x_{1},\ldots,x_{n}\}$, consider the augmented configuration matrix $\mathcal{X}$ introduced in section 4. If $W=(w_{1},\ldots,w_{n})\in\mathbb{W}_{0}(X)=\text{Ker}\,\mathcal{X}$, that is, the corresponding weighted system $(X,w)$ has $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{O}$ then, according to corollary 2.6, $\mu_{2}=-w_{0}$ for some constant $w_{0}\in\mathbb{R}$. In particular, it follows that
(A.2.1) $\mu_{2}(x_{i})=\sum_{j=1}^{n}w_{j}s_{ij}=-w_{0},\quad i=1,\ldots,n.$
Moreover, if $X$ is co-spherical, i.e., if there is a point $p$ equidistant from all points in $X$, and if $R^{2}=\overrightarrow{px_{i}}^{2}$, then $\mu_{2}(p)=\mu_{0}R^{2}=0$. So $w_{0}=0$ and, by equation (A.2.1), $W\in\text{Ker}\,\mathcal{B}(X)$, where
(A.2.2) $\mathcal{B}(X)=\left[\begin{array}[]{cccc}0&s_{12}&\cdots&s_{1n}\\ s_{21}&0&\cdots&s_{2n}\\ \vdots&\vdots&\ddots&\vdots\\ s_{n1}&s_{n2}&\cdots&0\end{array}\right].$
If $X$ has codimension $c>0$, the above implies that the determinant of $\mathcal{B}(X)$ is zero, which is the condition found in [10]. Cayley observed that if $n=4$ and $c=1$, a factorization of $\det\mathcal{B}(X)$ produces the classical Ptolemy identity for four cocircular points.
Let us view $\mu_{0}$ as a linear functional on the weight space $\mathbb{W}$ introduced in section 4. Notice corollary 2.6 ensures that $\text{Ker}\,\mathcal{B}(X)\cap\mu_{0}^{-1}(0)\leq\mathbb{W}_{0}(X)$ for arbitrary $X$, with equality occurring if and only if $\mu_{2}=0$ for any basis of $\mathbb{W}_{0}(X)$, a condition which is fulfilled, as we just saw, if $X$ is co-spherical. It is noteworthy that the matrix $\mathcal{B}(X)$ plays a key role in the reduction of the $n$-body problem by rotations and translations [3, 28]. Indeed, up to a multiplicative constant, $\mathcal{B}(X)$ represents a bilinear form known as the relative configuration matrix of $X$, for it actually characterizes $X$ up to isometries [3, 28]. We shall consider the kernel of $\mathcal{B}(X)$, viewed as a bilinear form, in appendix B.
Now consider the matrix
(A.2.3) $\mathcal{C}(X)=\left[\begin{array}[]{cc}0&\bm{1}^{T}\\ \bm{1}&\mathcal{B}(X)\end{array}\right],$
where $\bm{1}$ is the $n\times 1$ matrix $1\cdots 1]^{T}$. We see from equation (A.2.1) that $W\in\mathbb{W}_{0}(X)$ if and only if $\widetilde{W}=(w_{0},w_{1},\ldots,w_{n})\in\text{Ker}\,\,\mathcal{C}(X)$, with $w_{0}=-\sum_{j=1}^{n}w_{j}\overrightarrow{x_{i}x_{j}}^{2}$, for any $i$. The determinant of $\mathcal{C}(X)$ is known as the Cayley-Menger determinant of $X$ [8, 9].
The preceding paragraphs contain a proof of the statement below.
Theorem A.2.1.
For any configuration $X$ we have that $\emph{Ker}\,\,\mathcal{B}(X)\cap\mu_{0}^{-1}(0)\leq\mathbb{W}_{0}(X)\cong\emph{Ker}\,\,\mathcal{C}(X)$, with $\emph{Ker}\,\,\mathcal{B}(X)\cap\mu_{0}^{-1}(0)=\mathbb{W}_{0}(X)$ if and only if for every weighted system $(X,w)$ with $\mu_{0}=0$ and $\mu_{1}=\overrightarrow{O}$, we also have $\mu_{2}=0$. The latter conditions are fulfilled if $X$ is a co-spherical configuration.
The aforementioned natural isomorphism $W\mapsto\widetilde{W}$ in theorem A.2.1 appeared in corollary 3.4 of [14], where the conclusions in the next paragraph can also be found.
Theorem A.2.1 implies that $\mathcal{C}(X)$ has nullity equal to $\dim\mathbb{W}_{0}(X)$, which is $c$, by lemma 4.1. For $c=0$, the rank of $\mathcal{C}(X)$ is thus $n+1$, so $\det\mathcal{C}(X)\neq 0$. This means that there are no constraints on mutual distances, as expected. For configurations of codimension $c>0$, $\mathcal{C}(X)$ has rank nc+1n-c+1, so all subdeterminants of $\det\mathcal{C}(X)$ of size $(n-c+2)\times(n-c+2)$ must vanish. In particular, the $(n-c+2)\times(n-c+2)$ principal minors of $\mathcal{C}(X)$, which are precisely the Cayley-Menger determinants of the Dziobek subconfigurations of $X$ considered in section 4, are all equal to zero. We shall soon check that not all Cayley-Menger determinants of Dziobek subconfigurations are independent and that, in general, a much smaller number of such determinants suffices to guarantee the correct codimension for $X$.
Let us compute the number of constraints that must be satisfied by the mutual distances of a configuration of $n$ points and dimension $d$. Let $E(d)$ be the group of Euclidean (rigid) motions in $\mathbb{R}^{d}$. The manifold of congruence classes of configurations of $n$ points in $\mathbb{R}^{d}$ modulo isometries is $\mathbb{R}^{dn}\big/E(d)$, which has dimension dnd(d+1)/2dn-d(d+1)/2. The natural coordinates to describe the congruence classes are the $n\choose 2$ mutual distances between the points. For configurations of codimension c=(n1)d>0c=(n-1)-d>0, the number of mutual distances exceeds the dimension of the manifold of congruence classes by
(A.2.4) ${n\choose 2}-\left[dn-\frac{d(d+1)}{2}\right]={c+1\choose 2}$,
which is the number of constraints that must be imposed on the mutual distances.
Theorem A.2.2.
Let $X$ be a configuration of dimension d<n1d<n-1. The mutual distances of $X$ are such that $c+1\choose 2$} functionally independent Cayley-Menger determinants of subconfigurations of $X$ vanish, where $c=\emph{codim}(X)$.
Proof.
Inspired by the argument in section 3.5 of [21], we use induction on the codimension of $X$. When $c=1$, there is the unique constraint $\det\mathcal{C}(X)=0$, so the formula is valid in this case. Suppose the formula holds for a configuration $X$ of $n$ points and dimension $d$, where d=n(c+1)d=n-(c+1). Adding the point $p$ to $X$ without changing the dimension, the new configuration will have codimension n+1(d+1)=nd=c+1n+1-(d+1)=n-d=c+1. The additional constraints are the Cayley-Menger determinants of subconfigurations of $X\cup\{p\}$ formed by a selection of (any) $d$ points in $X$, the point $p$, and each of the remaining ndn-d points of $X$, one at a time. Thus we get a total of
${c+1\choose 2}+n-d={c+2\choose 2}$
constraints, as claimed. ∎
In the proof of theorem A.2.2, notice that it is possible to choose the subconfiguration of $X\cup\{p\}$ formed by the $d$ points of $X$ and $p$ as a core of $X\cup\{p\}$, cf. section 4. In this case, the additional contraints are expressed as Cayley-Menger determinants of Dziobek subconfigurations of $X\cup\{p\}$.
Appendix B Extended Leibniz Identities
In this appendix we show that the kernel of the relative configuration matrix $\mathcal{B}(X)$, viewed as a symmetric bilinear form on the space $\mathbb{D}=\mu_{0}^{-1}(0)$ is precisely $\mathbb{W}_{0}(X)$, the space of weight vectors with $\mu_{1}=\overrightarrow{O}$. This basic remark, which appears in [3], produces relations of interest to applications. Here we provide an independent proof, followed by an application to central configurations.
Firstly observe that if $W_{1}\in\mathbb{D}$ is in the kernel of $\mathcal{B}(X)$, then $W_{1}\in\mathbb{W}_{0}(X)$ by theorem A.2.1. Now let $W_{1}\in\mathbb{W}_{0}(X)$. According to theorem A.2.1, $\widetilde{W_{1}}=(w_{0},W_{1})\in\text{Ker}\,\,\mathcal{C}(X)$, thus, going back to equations (A.2.2) and (A.2.3), we have that
$\bm{1}^{T}W_{1}=0,\qquad w_{0}\bm{1}+\mathcal{B}(X)W_{1}=\bm{0},$
where $\bm{0}$ is the $n\times 1$ matrix $0\ldots 0]^{T}$. Let $W_{2}\in\mathbb{D}$. Multiplying both sides of the latter equation on the left by $W_{2}^{T}$ and using that $W_{2}^{T}\bm{1}=\bm{1}^{T}W_{2}=0$, we arrive at the extended Leibniz identity
$W_{2}^{T}\mathcal{B}(X)W_{1}=0,$
so the symmetric bilinear form defined by $\mathcal{B}(X)$ on the subspace $\mathbb{D}$ vanishes identically on $\mathbb{W}_{0}(X)$. Going back to equations (5.4) and theorem 5.2, we have the following corollary.
Corollary B.1.
A configuration of particles $(X,m)$ with $\mu_{0}\neq 0$ is a central configuration if and only if
$\sum_{i<k}m_{i}m_{k}S_{ij}S_{kl}s_{ik}=0,\quad j,l=1,\cdots,n.$
Acknowledgements. I wish to thank Alain Albouy for early discussions on the project that lead to this work, and for the comments and suggestions received from the participants of events organized by Luis Fernando Mello, Antonio Carlos Fernandes and Lei Zhao, and the Mathematical Physics Seminar in Recife. I specially thank Samuel Adrian Antz for sending a list of corrections to a previous version of the manuscript.